In order to economize on the oxygen supply in the space ships it has been suggested that the oxygen in exhaled CO2 be converted to water by reduction with hydrogen the CO2 output per astronaut has been estimated as one kg per 24 hours day and experimental catalytic converter reduces CO2 at a rate of 600 ml STP per minute what fraction of the time would such a converter have to operate in order to keep up the CO2 output of one astronaut

Grade:11

1 Answers

Avni Chauhan
5 years ago
moles of CO2= 1000/44
Volume = moles x 22.4L = 509 L
time required = total volume / rate
= 509000/600 = 848.48 min = 14.14 hrs = 58.9% of the day